# From Lengthy to Lucid: A Systematic Literature Review on NLP Techniques for Taming Long Sentences

### v1

@@ 3.5. Other Trends
In this subsection, we explore additional trends within the literature of sentence compression such as methods employed for length control as well as approaches related to multilingual sentence compression.
3.5.1. Length Control in Sentence Compression
@@
Length-controllable approaches for sentence summarization (Baziotis et al., 2019; Zhou and Rush, 2019) are attracting increasing attention during the last years. Several studies incorporate an implicit length control to improve the quality of the generated compressions, while others control the length explicitly during inference. For example, (Zhou and Rush, 2019) applies a length penalty to avoid very short compressions, while explicit approaches use decoder embeddings (Févry and Phang, 2018; Baziotis et al., 2019), discrete optimization search (Schumann et al., 2020; Liu et al., 2022a) or dynamic programming (Liu et al., 2022b). This subsection discusses explicit control in sentence summarization.
@@
Explicit length-controllable models can be either character-level (Napoles et al., 2011a; Liu et al., 2022b; Sclar et al., 2022) or word-level (Févry and Phang, 2018; Baziotis et al., 2019; Schumann et al., 2020; Liu et al., 2022a). The first abstractive work to perform length control at the character level is introduced in (Napoles et al., 2011a), where the length control is achieved based on dynamic programming to find the optimal number of paraphrases for reducing the character length of the compression. Recently, another character-length control approach was proposed in (Liu et al., 2022b), where the length control is formalized as a 0/1 knapsack problem, where the value is formulated as the predicted word’s log-likelihood and the weight is represented as the number of characters in the word. For more efficient inference, different lengths are divided into buckets. This method can be either supervised (using existing sentence compression datasets) or unsupervised (using pseudo-compressions generated by (Schumann et al., 2020)). Finally, REFEREE (Sclar et al., 2022) is a sentence compression method that is based on symbolic knowledge distillation (West et al., 2022), where knowledge is transferred from a large language model to a significantly smaller one. REFEREE is composed of two models: a distillation model (REFEREE-distill) which transfers the knowledge to the smaller model, and a length-controllable model (REFEREE-control) which is trained on the generated examples from the distillation process.
@@
Despite some efforts for character-level length control, the majority of the existing literature focuses on word-level length control in an unsupervised setting. More specifically, Schumann et al. (2020) propose an unsupervised length-controllable method for extractive sentence compression that is based on a discrete search optimization strategy. The length control is performed at the word level by extracting source words from the input sentence based on hard length constraints. The best sentence compression is selected based on a first-choice hill-climbing (Raffel et al., 2017) approach according to a hand-crafted heuristic function that aims to maximize predefined aspects such as fluency and similarity to the source. This approach can explicitly control the length of the generated compressions by searching across the space of all the feasible compressions. Despite the high performance of this method, it is slow and computationally exhausting. Another unsupervised word-level length controllable approach is proposed in (Liu et al., 2022a) where dynamic programming decoding is employed to control the length of the output. This approach allows for faster inference compared to (Schumann et al., 2020), since the exhaustive search is performed only for generating the data for the model’s training. Both methods are extractive since they rely only on deletion operations. On the other hand, existing auto-encoder approaches can be either extractive (Févry and Phang, 2018) or abstractive (Baziotis et al., 2019). The length control is achieved in both approaches with a length embedding added to the decoder to influence the length of the generated compressions.
3.5.2. Multilingual Sentence Compression
@@
Besides monolingual sentence compression, there have been a few attempts at multilingual (Rodrigues et al., 2018) and cross-lingual sentence compression (Aziz et al., 2012; Duan et al., 2019b). In multilingual sentence compression, the goal is to generate compressed versions of a given sentence in two or more different languages, while in cross-lingual sentence compression, the target compression should be in a different language from the input sentence. A multilingual approach for both English and Portuguese was proposed in (Rodrigues et al., 2018), where multilingual representations were obtained by mapping monolingual embeddings into a common multilingual vector space. On the other hand, cross-lingual techniques work either by translating the output to the target language (Aziz et al., 2012) or by performing zero-shot techniques without the need for annotated data (Duan et al., 2019a). Even though the majority of monolingual works study the task for the English language, there is also literature for Chinese (Xu and Grishman, 2009; Zhao et al., 2022; Zi et al., 2021; Zhang et al., 2013b, a), Japanese (Hasegawa et al., 2017; Hirao et al., 2009; Harashima and Kurohashi, 2012; Oguro et al., 2002) and Portuguese (Nóbrega et al., 2020; Rodrigues et al., 2018) as well.
4. Sentence Splitting
@@ 4.1. Methods
The importance of sentence splitting lies in its ability to break down complex piece of text into shorter and coherent segments. There have been efforts (Nomoto, 2022) to analyze whether sentence splitting enhances the readability of a complicated sentence. In this context, annotators are asked to rate two-sentence and three-sentence splits from existing splitting datasets. Then, a Bayesian logistic regression model is used to examine the ratings. This model employs different linguistic and cognitive features, including cohesion, perception, and structure, among others. The primary finding of this study indicates that sentence splitting indeed allows for a better understanding of the text. In addition, two-sentence splits are preferred compared to three-sentence ones.
@@
In the majority of existing works, sentence splitting is typically met as an intermediate component of sentence simplification or discourse relation systems (Yu et al., 2019), rather than being handled as an independent task. For example, as part of the DISRPT 2019 Shared Task for identifying discourse relations in text spans, Yu et al. (2019) employ a sentence splitter using a stacking ensemble method leveraging the outputs of multiple models including deep learning, machine learning, and rule-based techniques. In addition, sentence splitting is also used as a preprocessing step of machine translation methods (Lee et al., 2008).
@@
Methods for sentence splitting can be syntax-based (Siddharthan, 2006; Siddharthan and Angrosh, 2014), discourse-based (Cripwell et al., 2021) or semantic-based (Kim et al., 2021). Syntax-based methods employ syntax or grammar rules to split the input sentence into shorter ones, while discourse-based methods consider the discourse relations within the input sentence. On the other hand, semantic-based methods typically rely on models trained on existing splitting corpora to learn to split the input sentence according to its semantics.
@@
Figure 5 shows the number of publications from 2005 to 2022 along with landmarks of key technology adoption. A significant increase in interest is evident after 2017 possibly driven by the introduction of DL methods such as RNNs and Transformers. This growth further highlights the importance of sentence splitting for improving various NLP tasks.
@@
Figure 5. Moving average of publications per year from 2005 to 2022 for sentence splitting along with landmarks of key technology adoption.
4.1.1. Syntax-based methods
@@
First works on sentence splitting appear to be closely related to sentence simplification (Siddharthan, 2006; Siddharthan and Angrosh, 2014). These works formulate the task of sentence splitting as a constraint satisfaction problem where the goal is to find a valid solution that satisfies a set of constraints. For example, Siddharthan (2006) creates a set of manually written grammar and syntactic constraints that were then enriched to form a larger set of more extensive ones (Siddharthan and Angrosh, 2014)Lee et al. (2008) start with an initial set of splitting rules and then use a transformation-based learning method to iteratively refine them. Lee and Don (2017) employ a decision tree with different attributes such as sentence length, syntactic dependency patterns and POS labeling.
4.1.2. Discourse-based methods
@@
Cripwell et al. (2021) introduce a sentence-splitting model based on discourse structure. This method suggests two approaches: a discourse tree model as well a Transformer model. As a result, it falls both into discourse-based and semantic-based categories. This approach ensures that the discourse relation in the original long sentence is preserved in the output of shortened sentences, but requires the presence of a discourse connective, which is not always present when dealing with real sentences.
4.1.3. Semantic-based methods
@@
Recent work addresses the task of sentence splitting with semantic-based methods using supervised training of sequence-to-sequence models on splitting datasets such as WebSplit (Narayan et al., 2017; Aharoni and Goldberg, 2018), WikiSplit (Botha et al., 2018) and BiSECT (Kim et al., 2021). For example, Narayan et al. (2017) use a three-layered LSTM, while other approaches (Aharoni and Goldberg, 2018; Botha et al., 2018; Zhang et al., 2020b) use a bi-directional attention-based LSTM (Bahdanau et al., 2014) with a copying mechanism similar to a pointer-generator network (See et al., 2017). In (Wang et al., 2020) a separator network is used to enhance the standard attention-based LSTM. The separator network determines the content of each split and then forms a representation that is passed to the decoder for generating the split sentence. This method mitigates the repetition issues of previous models that are typically found within the split sentences. Another method to address repetition issues is proposed in (Fan et al., 2020), where a memory-based Transformer model is introduced. This method incorporates a memory gate layer positioned between the encoder and the decoder to filter out previously decoded information. In addition, the model is trained with a multi-task learning objective that exploits (Resource Description Framework) RDF triples to encourage the generation of factually correct sentences.
@@
Berlanga Neto and Ruiz (2021) adopt a cross-lingual approach for both English and Portuguese. They use a GRU model with an attention mechanism that is trained with a symbolic vocabulary. The symbolic vocabulary, which is based on POS tags, allows for generalization in different languages. To transform the generated symbolic tokens into actual text, different versions of MLM BERT models are used.  Kim et al. (2021) employ the Transformer architecture to propose several baseline models. More specifically, this works identifies three types of splits according to the volume of modifications: a) direct insertion, b) changes near split, and c) changes across the sentence. Direction insertion involves minimal or no modifications of the initial sentence while changes near split require slight adjustments around the splitting point. Changes across sentence involve significant changes to reconstruct the initial sentence. This works employs a BERT model with an custom loss function that takes into account these split categories. Guo et al. (2020) propose a multi-task learning approach to learn sentence splitting and factual correctness. A sequence-to-sequence model is trained not only for sentence splitting but also for determining the correctness of the given facts according to the input sentence. This approach is applied to both LSTM and Transformer models.
@@
Table 5 provides a list of splitting methods along with their corresponding category (semantic, discourse, or syntax-based). While earlier approaches mainly rely on syntax-based approaches that formulate the task with constraints (Siddharthan, 2006; Siddharthan and Angrosh, 2014) or employ simple rule-based methods (Lee et al., 2008), the growth of DL methods (RNNs and Transformers) led to the rise of semantic-based approaches (Narayan et al., 2017; Aharoni and Goldberg, 2018; Kim et al., 2021).
@@
Table 5. Overview of splitting methods based on category (syntax, discourse and semantic-based) and employed approach (constraints, decision trees, discourse trees, RNNs and Transformers). All methods except constraint-based methods are supervised.
Paper
Category Approach
@@
(Narayan et al., 2017; Aharoni and Goldberg, 2018; Botha et al., 2018; Zhang et al., 2020b; Berlanga Neto and Ruiz, 2021)
(Fan et al., 2020; Kim et al., 2021)
(Guo et al., 2020)
(Siddharthan, 2006; Siddharthan and Angrosh, 2014)
(Lee et al., 2008)
(Lee and Don, 2017)
(Cripwell et al., 2021)
4.2. Datasets
@@
WebSplit (Narayan et al., 2017) is a large-scale synthetic dataset that consists of 5,546 complex sentences extracted by the RDF triples of the WebNLG (Gardent et al., 2017) corpus. The WebNLG corpus consists of RDF triples along with one or more verbalizations, i.e., RDF triples converted to human-readable text. The RDF triples contain relations between the different entities for each text. The process of constructing the WebSplit dataset begins with an initial sentence and its associated RDF triples. A subset of RDF triples for this sentence is identified, and the verbalizations of these RDF triples (shorter sentences) are then merged to create the split version of the original input sentence.
@@
Zhang et al. (2020b) reveal that WebSplit follows easily exploitable syntactic patterns that even simple rule-based techniques can recognize. An improved version of WebSplit is proposed in (Aharoni and Goldberg, 2018) that introduced an alternative training, validation and test set split with less overlap between splits. Still, WebSplit uses a relatively small vocabulary (3,311 words) and contains some unnatural expressions due to its synthetic nature.
@@
WikiSplit (Botha et al., 2018) contains one million natural English sentence splits that are created automatically from user edits of Wikipedia revision history. WikiSplit’s vocabulary is 90 times larger than WebSplit. At the same time it provides more coherent and naturally occurring examples. However, since the dataset is constructed automatically, the preservation of meaning within the examples cannot be guaranteed. As a result, this dataset is typically used to augment training data for enhancing the performance of splitting models.
@@
Wiki-BM (Zhang et al., 2020b) is a subset of the original test set of WikiSplit that contains 500 human-annotated examples out of the total 5,000. MinWikiSplit (Niklaus et al., 2019) is a large-scale dataset for sentence splitting that builds upon WikiSplit using the principle of minimal prepositions. It contains 203K complex sentences paired with their split sentences according to the principle of minimality, where each sentence is split into a set of minimal propositions. It was originally proposed to improve WikiSplit’s split sentences since some sentences could be further split. Each sentence in MinWikiSplit is split into minimally shortened sentences based on human-written, handcrafted rules.
@@
BiSECT (Kim et al., 2021) contains one million complex sentences accompanied with shorter sentences in four different languages including English, German, French and Spanish. The dataset is created with bilingual pivoting by mapping 1-2 and 2-1 sentence pairs in a bilingual dataset and then translating the pairs into one language. Even though this dataset is typically used for sentence simplification, in this study we focus on the reduction of the sentence length without taking the content into account.
Table 6. Datasets for sentence splitting.
@@
WebSplit (v.0.1) (Narayan et al., 2017) 5,546 2017
WebSplit (v.1.0) (Aharoni and Goldberg, 2018) 18,830 2018
WikiSplit (Botha et al., 2018) 999,944 2018
MinWikiSplit (Niklaus et al., 2019) 203K 2019
Wiki-BM (Zhang et al., 2020b) 500 2020
BiSECT (Kim et al., 2021) 1M 2021
4.3. Evaluation Metrics
The most commonly used metric for assessing the performance of sentence splitting models is BLEU (Papineni et al., 2002), which is based on n-gram precision. BLEU is used to compute the similarity between a ground truth set of sentences and a predicted set of sentences as follows:
(2) $\textrm{BLEU}=\textrm{BP}\cdot\exp\bigg{(}\sum_{n=1}^{N}w_{n}\log p_{n}\bigg{)},$

### v2

@@ 3.5 Other Trends
In this subsection, we explore additional trends within the literature of sentence compression such as methods employed for length control as well as approaches related to multilingual and multimodal sentence compression.
3.5.1 Length Control in Sentence Compression
@@
Length-controllable approaches for sentence summarization Baziotis2019 ; zhou2019 are attracting increasing attention during the last years. Several studies incorporate an implicit length control to improve the quality of the generated compressions, while others control the length explicitly during inference. For example, zhou2019 applies a length penalty to avoid very short compressions, while explicit approaches use decoder embeddings fevry2018unsupervised ; Baziotis2019 , discrete optimization search Schumann2020DiscreteExtraction ; liu2022learning or dynamic programming Liu2022Controlling . This subsection discusses explicit control in sentence summarization.
@@
Explicit length-controllable models can be either character-level napoles2011paraphrastic ; Liu2022Controlling ; SclarReferee2022 or word-level fevry2018unsupervised ; Baziotis2019 ; Schumann2020DiscreteExtraction ; liu2022learning ; juseon2024instructcmp . The first abstractive work to perform length control at the character level is introduced in napoles2011paraphrastic , where the length control is achieved based on dynamic programming to find the optimal number of paraphrases for reducing the character length of the compression. Recently, another character-length control approach was proposed in Liu2022Controlling , where the length control is formalized as a 0/1 knapsack problem, where the value is formulated as the predicted word’s log-likelihood and the weight is represented as the number of characters in the word. For more efficient inference, different lengths are divided into buckets. This method can be either supervised (using existing sentence compression datasets) or unsupervised (using pseudo-compressions generated by Schumann2020DiscreteExtraction ). Finally, REFEREE SclarReferee2022 is a sentence compression method that is based on symbolic knowledge distillation west2022symbolic , where knowledge is transferred from a large language model to a significantly smaller one. REFEREE is composed of two models: a distillation model (REFEREE-distill) which transfers the knowledge to the smaller model, and a length-controllable model (REFEREE-control) which is trained on the generated examples from the distillation process.
@@
Despite some efforts for character-level length control, the majority of the existing literature focuses on word-level length control in an unsupervised setting. More specifically, Schumann2020DiscreteExtraction propose an unsupervised length-controllable method for extractive sentence compression that is based on a discrete search optimization strategy. The length control is performed at the word level by extracting source words from the input sentence based on hard length constraints. The best sentence compression is selected based on a first-choice hill-climbing raffel17a approach according to a hand-crafted heuristic function that aims to maximize predefined aspects such as fluency and similarity to the source. This approach can explicitly control the length of the generated compressions by searching across the space of all the feasible compressions. Despite the high performance of this method, it is slow and computationally exhausting. Another unsupervised word-level length controllable approach is proposed in liu2022learning where dynamic programming decoding is employed to control the length of the output. This approach allows for faster inference compared to Schumann2020DiscreteExtraction , since the exhaustive search is performed only for generating the data for the model’s training. Both methods are extractive since they rely only on deletion operations. On the other hand, existing auto-encoder approaches can be either extractive fevry2018unsupervised or abstractive Baziotis2019 . The length control is achieved in both approaches with a length embedding added to the decoder to influence the length of the generated compressions. Finally, InstructCMP juseon2024instructcmp introduces a zero-shot length control approach that employs Large Language Models (LLMs), such as ChatGPT, to restrict the length of the compression. The best results are obtained when the prompt instructions include both the desired target length of the compression and the original sentence’s word count, along with the number of words to remove. Finally, a recent line of work introduces explicit length control for cross-lingual compression, where the desired compression ratio is achieved through a token prepended to the input, enabling a single model to generate translations at different target lengths tollef2025cross .
3.5.2 Multilingual Sentence Compression
@@
Besides monolingual sentence compression, there have been a few attempts at multilingual frederico2018multilingual and cross-lingual sentence compression aziz2012cross ; duan2019contrastive ; gasan2024multi ; tollef2025cross . In multilingual sentence compression, the goal is to generate compressed versions of a given sentence in two or more different languages, while in cross-lingual sentence compression, the target compression should be in a different language from the input sentence. A multilingual approach for both English and Portuguese was proposed in frederico2018multilingual , where multilingual representations were obtained by mapping monolingual embeddings into a common multilingual vector space. On the other hand, cross-lingual techniques work either by translating the output to the target language aziz2012cross ; gasan2024multi , jointly learning translation and compression from compressed parallel corpora tollef2025cross , or performing zero-shot techniques without the need for annotated data duan2019zero . Even though the majority of monolingual works study the task for the English language, there is also literature for Chinese xu2009parse ; zhao2022simple ; zi2021som ; zhang2012research ; zhang2013chinese , Japanese hasegawa2017japanese ; hirao2009syntax ; harashima2012flexible ; oguro2002evaluation , Portuguese nobrega2020sentence ; frederico2018multilingual , and Persian tavakoli2025sentence as well.
3.5.3 Multimodal Sentence Compression
Multimodal sentence summarization (MMSS), a relatively new field in the era of multimodality, refers to the process of generating short and informative summaries by integrating information from multiple data types such as text, images, audio and videos. Unlike traditional sentence summarization, which relies only on text, multimodal summarization combines these sources to provide a summary that reflects information from the different content types.
Existing approaches mainly focus on generating compressions by combining information from a source image and text. For example, BART-MMSS Lin2023Multimodal adapts BART for multimodal sentence summarization by integrating image information using a Vision Transformer, which extracts image embeddings and injects them into the BART model. In addition, it employs an explicit critical token learning module to identify important tokens in the source sentence, aligning them with visual information in the image, to ensure that they will be included in the final summary.
Another approach that employs generative pre-trained models for MMSS is Vision-GPLM jing2023vision . Vision-GPLM uses two different encoders for text and image features and merges them through multi-head attention. The model is trained in two phases: a pre-training phase, where only the visual encoder is trained, and a fine-tuning phase, where all components are trained jointly for MMSS.
yuan2024exploring introduce a framework based on the Information Bottleneck principle to more effectively manage the incorporation of visual elements. While existing methods often risk including too much (over-preservation) or too little (over-compression) visual detail Lin2023Multimodal , this approach works by creating a maximally compressed mapping of the input data, where only information essential to the summary is retained.
4 Sentence Splitting
@@ 4.1 Methods
The importance of sentence splitting lies in its ability to break down complex piece of text into shorter and coherent segments. There have been efforts nomoto2022fewer to analyze whether sentence splitting enhances the readability of a complicated sentence. In this context, annotators are asked to rate two-sentence and three-sentence splits from existing splitting datasets. Then, a Bayesian logistic regression model is used to examine the ratings. This model employs different linguistic and cognitive features, including cohesion, perception, and structure, among others. The primary finding of this study indicates that sentence splitting indeed allows for a better understanding of the text. In addition, readability is found to improve more with a two-sentence split than with additional splits nomoto2023does .
@@
In the majority of existing works, sentence splitting is typically met as an intermediate component of sentence simplification liu2025community or discourse relation systems yu2019gumdrop , rather than being handled as an independent task. For example, as part of the DISRPT 2019 Shared Task for identifying discourse relations in text spans, yu2019gumdrop employ a sentence splitter using a stacking ensemble method leveraging the outputs of multiple models including deep learning, machine learning, and rule-based techniques. In addition, sentence splitting is also used as a preprocessing step of machine translation methods lee2008transformation .
@@
Methods for sentence splitting can be syntax-based Siddharthan2006 ; Siddharthan2014 , discourse-based Cripwell2021 or semantic-based kim-etal-2021-bisect . Syntax-based methods employ syntax or grammar rules to split the input sentence into shorter ones, while discourse-based methods consider the discourse relations within the input sentence. On the other hand, semantic-based methods typically rely on models trained on existing splitting corpora to learn to split the input sentence according to its semantics.
@@
Figure 5 shows the cumulative number of publications from 2005 to 2024 along with landmarks of key technology adoption. A significant increase in interest is evident after 2017 possibly driven by the introduction of DL methods such as RNNs and Transformers. This growth further highlights the importance of sentence splitting for improving various NLP tasks.
@@
Figure 5: Cumulative number of publications per year from 2005 to 2025 for sentence splitting along with landmarks of key technology adoption.
4.1.1 Syntax-based methods
@@
First works on sentence splitting appear to be closely related to sentence simplification Siddharthan2006 ; Siddharthan2014 . These works formulate the task of sentence splitting as a constraint satisfaction problem where the goal is to find a valid solution that satisfies a set of constraints. For example, Siddharthan2006 creates a set of manually written grammar and syntactic constraints that were then enriched to form a larger set of more extensive ones Siddharthan2014 lee2008transformation start with an initial set of splitting rules and then use a transformation-based learning method to iteratively refine them. lee2017splitting employ a decision tree with different attributes such as sentence length, syntactic dependency patterns and POS labeling.
4.1.2 Discourse-based methods
@@
Cripwell2021 introduce a sentence-splitting model based on discourse structure. This method suggests two approaches: a discourse tree model as well a Transformer model. As a result, it falls both into discourse-based and semantic-based categories. This approach ensures that the discourse relation in the original long sentence is preserved in the output of shortened sentences, but requires the presence of a discourse connective, which is not always present when dealing with real sentences.
4.1.3 Semantic-based methods
@@
Recent work addresses the task of sentence splitting with semantic-based methods using supervised training of sequence-to-sequence models on splitting datasets such as WebSplit narayan-etal-2017-split ; aharoni-goldberg-2018-split , WikiSplit botha-etal-2018-learning and BiSECT kim-etal-2021-bisect . For example, narayan-etal-2017-split use a three-layered LSTM, while other approaches aharoni-goldberg-2018-split ; botha-etal-2018-learning ; Zhang2020 use a bi-directional attention-based LSTM Bahdanau2014 with a copying mechanism similar to a pointer-generator network See2017GetNetworks . In wang2020hierarchical a separator network is used to enhance the standard attention-based LSTM. The separator network determines the content of each split and then forms a representation that is passed to the decoder for generating the split sentence. This method mitigates the repetition issues of previous models that are typically found within the split sentences. Another method to address repetition issues is proposed in fan2020memory , where a memory-based Transformer model is introduced. This method incorporates a memory gate layer positioned between the encoder and the decoder to filter out previously decoded information. In addition, the model is trained with a multi-task learning objective that exploits (Resource Description Framework) RDF triples to encourage the generation of factually correct sentences.
@@
berlanga2021split adopt a cross-lingual approach for both English and Portuguese. They use a GRU model with an attention mechanism that is trained with a symbolic vocabulary. The symbolic vocabulary, which is based on POS tags, allows for generalization in different languages. To transform the generated symbolic tokens into actual text, different versions of MLM BERT models are used.  kim-etal-2021-bisect employ the Transformer architecture to propose several baseline models. More specifically, this works identifies three types of splits according to the volume of modifications: a) direct insertion, b) changes near split, and c) changes across the sentence. Direction insertion involves minimal or no modifications of the initial sentence while changes near split require slight adjustments around the splitting point. Changes across sentence involve significant changes to reconstruct the initial sentence. This works employs a BERT model with an custom loss function that takes into account these split categories. guo2020fact propose a multi-task learning approach to learn sentence splitting and factual correctness. A sequence-to-sequence model is trained not only for sentence splitting but also for determining the correctness of the given facts according to the input sentence. This approach is applied to both LSTM and Transformer models. Another Transformer-based method is introduced in alajlouni2023knowledge , where a split-and-rephrase approach is proposed for transferring knowledge from a rule-based model to a pre-trained language model. First, candidate splits are generated based on the rule-based model, which are then filtered according to meaning preservation, grammatical correctness and simplicity. The filtered samples are used to fine-tune a T5 Transformer model for sentence splitting. Although most research focuses on English, sentence splitting has also been examined in typologically diverse languages such as Persian talebpour2024split , where encoder–decoder models have shown strong performance, demonstrating the broader cross-lingual applicability of the task.
Even though LLMs have achieved great results in various NLP tasks, there is only one work that is focused on their use for sentence splitting ponce-etal-2024-split . More specifically, ponce-etal-2024-split evaluate different LLMs such as Alpaca, GPT-4 Turbo and Pythia, for the split and rephrase task. LLMs were prompted to generate split versions of complex sentences, in different settings, including zero-shot, few-shot in-context learning, and fine-tuning. Regarding fine-tuning, results show that around 3,000 examples of complex and split sentences are needed to achieve reasonable accuracy, with 4,000 to 6,000 pairs to yield the best performance depending on the domain. In addition, instruction-tuned models show promising results with as few as five curated examples.
@@
Table 5 provides a list of splitting methods along with their corresponding category (semantic, discourse, or syntax-based). While earlier approaches mainly rely on syntax-based approaches that formulate the task with constraints Siddharthan2006 ; Siddharthan2014 or employ simple rule-based methods lee2008transformation , the growth of DL methods (RNNs and Transformers) led to the rise of semantic-based approaches narayan-etal-2017-split ; aharoni-goldberg-2018-split ; kim-etal-2021-bisect .
@@
Table 5: Overview of splitting methods based on category (syntax, discourse and semantic-based) and employed approach (constraints, decision trees, discourse trees, RNNs and Transformers). All methods except constraint-based methods are supervised.
Paper Category Approach
@@
narayan-etal-2017-split ; aharoni-goldberg-2018-split ; botha-etal-2018-learning ; Zhang2020 ; berlanga2021split
fan2020memory ; kim-etal-2021-bisect ; ponce-etal-2024-split ; talebpour2024split
alajlouni2023knowledge
guo2020fact
Siddharthan2006 ; Siddharthan2014
lee2008transformation
lee2017splitting
Cripwell2021
4.2 Datasets
@@
WebSplit narayan-etal-2017-split is a large-scale synthetic dataset that consists of 5,546 complex sentences extracted by the RDF triples of the WebNLG Gardent2017 corpus. The WebNLG corpus consists of RDF triples along with one or more verbalizations, i.e., RDF triples converted to human-readable text. The RDF triples contain relations between the different entities for each text. The process of constructing the WebSplit dataset begins with an initial sentence and its associated RDF triples. A subset of RDF triples for this sentence is identified, and the verbalizations of these RDF triples (shorter sentences) are then merged to create the split version of the original input sentence.
@@
Zhang2020 reveal that WebSplit follows easily exploitable syntactic patterns that even simple rule-based techniques can recognize. An improved version of WebSplit is proposed in aharoni-goldberg-2018-split that introduced an alternative training, validation and test set split with less overlap between splits. Still, WebSplit uses a relatively small vocabulary (3,311 words) and contains some unnatural expressions due to its synthetic nature.
@@
WikiSplit botha-etal-2018-learning contains one million natural English sentence splits that are created automatically from user edits of Wikipedia revision history. WikiSplit’s vocabulary is 90 times larger than WebSplit. At the same time it provides more coherent and naturally occurring examples. However, since the dataset is constructed automatically, the preservation of meaning within the examples cannot be guaranteed. As a result, this dataset is typically used to augment training data for enhancing the performance of splitting models.
@@
Wiki-BM Zhang2020 is a subset of the original test set of WikiSplit that contains 500 human-annotated examples out of the total 5,000. MinWikiSplit niklaus-etal-2019-minwikisplit is a large-scale dataset for sentence splitting that builds upon WikiSplit using the principle of minimal prepositions. It contains 203K complex sentences paired with their split sentences according to the principle of minimality, where each sentence is split into a set of minimal propositions. It was originally proposed to improve WikiSplit’s split sentences since some sentences could be further split. Each sentence in MinWikiSplit is split into minimally shortened sentences based on human-written, handcrafted rules.
@@
BiSECT kim-etal-2021-bisect contains one million complex sentences accompanied with shorter sentences in four different languages including English, German, French and Spanish. The dataset is created with bilingual pivoting by mapping 1-2 and 2-1 sentence pairs in a bilingual dataset and then translating the pairs into one language. Even though this dataset is typically used for sentence simplification, in this study we focus on the reduction of the sentence length without taking the content into account.
WikiSplit++ tsukagoshi2024wikisplit contains over 600,000 examples of sentence splits from WikiSplit, addressing issues with factual consistency and coherence in the data. Unlike WikiSplit, which may include hallucinations or inaccuracies due to its automated construction, WikiSplit++ uses a natural language inference classifier to filter out unreliable examples, specifically removing those where the simplified sentence does not match the meaning of the original complex sentence. Then, sentence order is reversed in the output to encourage models to generate more diversified outputs rather than simply copying the word sequence of the original complex sentence.
Table 6: Datasets for sentence splitting.
@@
WebSplit (v.0.1) narayan-etal-2017-split 5,546 2017
WebSplit (v.1.0) aharoni-goldberg-2018-split 18,830 2018
WikiSplit botha-etal-2018-learning 999,944 2018
MinWikiSplit niklaus-etal-2019-minwikisplit 203K 2019
Wiki-BM Zhang2020 500 2020
BiSECT kim-etal-2021-bisect 1M 2021
WikiSplit++ tsukagoshi2024wikisplit 630,433 2024
4.3 Evaluation Metrics
The most commonly used metric for assessing the performance of sentence splitting models is BLEU Papineni2002 , which is based on n-gram precision. BLEU is used to compute the similarity between a ground truth set of sentences and a predicted set of sentences as follows:
$\textrm{BLEU}=\textrm{BP}\cdot\exp\bigg(\sum_{n=1}^{N}w_{n}\log p_{n}\bigg),$ (2)